OpenStudy (anonymous):

Let f(x)=x2-16. Find f1(x)

OpenStudy (anonymous):

you substitute 1 for x so 2(1) - 16 f(x) = -14

OpenStudy (anonymous):

i think lol

OpenStudy (abmon98):

y2=x+16 y=(x+16)^1/2

OpenStudy (anonymous):

oh did you mean x^2??

OpenStudy (abmon98):

no square root

OpenStudy (anonymous):

Yea

OpenStudy (lena772):

The whole procedure is wrong if x is supposed to be squared, just saying.